OpenStudy (anonymous):

HELP! A bakery sells pies and cakes. Each pie takes 20 minutes to prepare and 45 minutes to cook. Each cake takes 30 minutes to cook and a total of 40 minutes to mix and decorate. In a given day, the bakery has 16 hours of employee time for food preparation and 3 ovens available for 8 hours of each. The bakery makes a profit of $5.50 on each pie and $6.00 on each cake. a) Let x = the number of pies and y = the number of cakes the bakery makes in a given day. Translate a system of linear inequalities. c) Find the number of each dessert the bakery should make each day to max its profit.

OpenStudy (michele_laino):

please, I'm trying....

OpenStudy (michele_laino):

I think, that we have three ovens for 16 hours, so it is like to have one oven which is availabile for 16*3=48 hours, do you agree?

OpenStudy (michele_laino):

because three ovens are available contemporarily

OpenStudy (michele_laino):

so 48 hours are equivalent to to 48*60 minutes, then your first inequality is: \[65*x+70*y \le 48*60\] do you agree?

OpenStudy (michele_laino):

@djibben615

OpenStudy (michele_laino):

so, after a simple simplification, I rewrite the first inequality as below: \[13x+14y \le576\]

OpenStudy (anonymous):

gotcha! What would the second system be?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

Now we have to write the second inequality do you agree?

OpenStudy (michele_laino):

I think that if our bakery sells only cakes, then he makes the maximum gain. Now we have to calculate maximum number of cake that our bakery can sell

OpenStudy (michele_laino):

I think that we can write: \[70*y \le 48*60\] so: \[y \le \frac{ 48*60 }{ 70 }\le \frac{ 49*60 }{ 70 }=42\] so our bakery can sell at maximum 42 cakes and can gain at maximum 42*6.00=252$

OpenStudy (anonymous):

gotcha! this makes so much more sense now! thanks!

OpenStudy (michele_laino):

then our second inequality can be this: \[5.5*x+6*y \le 252\]

OpenStudy (anonymous):

gotcha that looks right now

OpenStudy (anonymous):

\[5.5x + 6y <= 252\]

OpenStudy (michele_laino):

second part is slightly more subtle

OpenStudy (anonymous):

What about the pies tho? Part c asks the number of both pies and cakes it'll make to max the profits?

OpenStudy (michele_laino):

because we have to maximize the subsequent quantity: 5.5x+6y

OpenStudy (anonymous):

\[5.5x + 6y \le 252\] this is the second inequality of the system correct?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

So these make up the system?: \[\left\{ 13x + 14 y \le 576 \right\} \left\{ 5.5 + 6y \le 252 \right\}\]

OpenStudy (michele_laino):

yes I think!

OpenStudy (anonymous):

Okay, so now to work on C?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

I'm not sure that system will work since the second inequality has both x and y in it too. We have to be able to substitute the second inequality into the first one.

OpenStudy (michele_laino):

I think that we have to use only the first inequality and graph it

OpenStudy (anonymous):

oh really? Okay

OpenStudy (michele_laino):

please note that, we got maximum of both x and y, when we have: \[13x+14y=576\] or, in other words, when: \[576-13x\] is divisible by 14

OpenStudy (michele_laino):

I'm trying to find solution...

OpenStudy (michele_laino):

If we proceed by attempts, we will get: x=12, y=30

OpenStudy (anonymous):

How would i graph that then?

OpenStudy (michele_laino):

please you have to graph this function: \[y=\frac{ 576-13x }{ 14 }\] or: \[y=\frac{ 288 }{ 7 }-\frac{ 13 }{ 14 }x\]

OpenStudy (anonymous):

gotcha! Now, how to get max profits for both pies and cakes?

OpenStudy (michele_laino):

I think setting x=12 and y=30

OpenStudy (anonymous):

so..... how would i figure out the total? :) thanks

OpenStudy (michele_laino):

total is: 5.5*12+6.0*30=246.00$